Solve the system of equations. Determine the number of solutions for each.
{1/3x+y=9
{-2x+2y=-6

Answer by edjones(8007)


2/3x+2y=18 Multiply 2 times the upper equation
-2x+2y=-6
----------
8/3x =24 Subtract the lower equation from the upper
x=9 Multply both sides times 3/8
(-2*9)+2y=-6 Substitute 9 for x in either of the 2 equations
2y=12 Add 18 to both sides
y=6 Divide both sides by 2.